OpenStudy (anonymous):

Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase had a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius.

OpenStudy (anonymous):

@omartfresh

OpenStudy (anonymous):

This one is hard but i think this would help "let the radius be r cm so the area of the base is πr^2 the area of the sleeve = 2πr(50) = 100π πr^2 + 100π = 3000 πr^2 + 100π - 3000 = 0 using the formula and my calculator I got r = (-100π ± √( (100π)^ - 4(π)(-3000) )/(2π) = appr 8.79 or a negative so the radius is appr 8.79 check: area of base = 242.1 area of sleeve = 2757.9 total = 3000 " by the way this is not my work heres the website http://www.jiskha.com/display.cgi?id=1366042639 hope this helped.